# Population demographics of white dwarf binaries with intermediate separations: Gaia constraints on post-AGB mass transfer

### v1

@@ 5.1 Orbital evolution under MT
We also alter the eccentricity evolution so that the final eccentricity, $e_{f}$, of the wide post-MT systems are related to their initial values, $e_{i}$, by $e_{f}=e_{i}^{4}+0.05$. A gaussian noise with a standard deviation of 0.05 is also added, and an upper bound of $e_{f}=1$ was imposed. This is an ad-hoc prescription which allows us to better reproduce the true distribution. However, simulations of CEE from Glanz & Perets (2021) found a correlation between initial and final eccentricities at the end of the dynamical phase which motivates a choice of this kind relating the two.
5.2 WD IFMR
Figure 11: Core mass as a function of the stellar radius on the AGB for various initial stellar masses. The thicker line weight highlights regions where the radius is greater than at any previous point in the evolution, which are used in the interpolation.
We also adopt a different WD IFMR which takes into account the interruption of the progenitor’s evolution due to the onset of binary interaction. On average, this should lead to lower WD masses compared to those predicted from single star evolution and ameliorate the discrepancy between the observed and simulated WD mass distributions (Section 4). We estimate WD mass as the he_core_mass from the MIST evolutionary tracks at a given giant radius and initial mass. For each system which interacts on the RGB or AGB, the appropriate track corresponding to the closest initial mass is chosen and the core mass at $r_{\rm L,peri}$ is set to be the final WD mass. Figure 11 plots the core mass vs. radius on the AGB for a range of initial masses. Note that during thermal pulses, the growth becomes non-monotonic. However, this is not problematic for our model, as we only consider regions where the radius is larger than at any earlier point in the evolution (otherwise interaction would have occurred earlier). Moreover, we see that the core mass can exceed $1.4\,M_{\odot}$ at the final stages of the AGB phase for progenitor masses $\sim 6-8\,M_{\odot}$. In this case, we return to using the Weidemann (2000) IFMR. This is reasonable as the core is defined in terms of the helium mass fraction and not all of this material ends up burning into carbon and oxygen. A very small fraction of systems are in this regime so the end results are insensitive to this choice.
For stars that interact on the AGB, our adopted IFMR leads to WD masses that are $\sim 10-20\%$ lower than those predicted by the Weidemann (2000) relation for progenitor masses between $\sim 1.5$ and $4.5\,M_{\odot}$. This corresponds to WD masses between $\sim 0.55-0.85\,M_{\odot}$. Outside this range, our IFMR can lead to higher WD masses. For comparison, Shahaf (2025) used a different method on a spectroscopic sample of AU-scale WD + red giant binaries in open clusters and found that these WDs were $\sim 20\%$ less massive than isolated WDs, which they attributed to binary interaction (also see Ironi et al. 2025).
Combined with the changes in the assumed orbital evolution due to MT described above (Section 5.1), we will refer to this modified version of the simulated population as our “fiducial” model. They key components of this model are illustrated in Figure 8. This new simulated population leads to a modified mock NSS catalog which we once again process through the Shahaf et al. (2024) pipeline.
5.3 An $\alpha$-formalism model
@@
The $\alpha$-formalism (Livio & Soker, 1988; de Kool, 1990) is often used to describe the orbital evolution resulting from CEE (e.g. Zorotovic et al., 2010; Davis et al., 2012; Scherbak & Fuller, 2023). Here, we test whether the $\alpha$-formalism can predict the properties of WD+MS binaries formed via interaction on the AGB. At its core, the formalism is a statement of energy conservation:
@@
where $E_{\rm bind}$ is the binding energy of the donor’s envelope, and the right-hand side is a fraction $\alpha$ of the orbital energy loss. $M_{i}$ is the initial mass of the WD progenitor (i.e. the donor). In the classic formulation, $E_{\rm bind}=-(GM_{i}M_{i,\rm env})/(\lambda R_{i})$ where $M_{i,\rm env}$ and $R_{i}$ are the envelope mass and radius of the donor, and $\lambda$ is a free parameter which encodes information about the envelope structure, typically set to $\sim 0.5$. Taking $R_{i}$ to be the Roche lobe radius at periastron and $M_{i,\rm env}=M_{i}-M_{\rm WD}$, we can solve for $a_{f}$ with one free parameter, namely the product $\alpha\lambda$.
@@
A value of $\alpha\lambda=1$ is typical under the standard assumptions because $\alpha$ is defined to be less than or equal to 1 and considering only the gravitational binding energy, we expect $\lambda$ to be on the order unity. A higher value may be understood as the donor having a particularly loosely bound envelope, which may be the case if it is undergoing a thermal pulse on the AGB, or if additional sources of energy (e.g. recombination) aid in the unbinding (e.g. Ivanova et al., 2013, 2015; Belloni et al., 2024; Yamaguchi et al., 2024a, a).
@@
We use the same WD IFMR as in Section 5.2 for this model, which we will refer to as the “$\alpha\lambda$ model” hereafter. Although the $\alpha$-formalism is not expected to provide an accurate description of the mass transfer process, we include this model as it may be useful to have an effective value of $\alpha\lambda$ to predict properties of PCEBs without having to use detailed evolutionary models, which themselves struggle to predict the outcome of CEE. The results of using this model can be found in Section 6.3.
@@ 6 Results of modified models
6.1 Fiducial model
The distributions of stellar and orbital parameters of the mock NCE sample from our fiducial model are plotted with a solid purple border on Figure 7.
The primary mass distribution of the systems in our fiducial model is more peaked compared to that of the E24 model, in better agreement with the observed distribution. This is the result of introducing a mass ratio dependence to the ratio between initial and final separation (Section 5.1). In this scheme, more massive MS companions are favored to form wide post-MT systems that are detectable by astrometry. Our model peaks at a slightly lower primary mass than the observed distribution. While this could be improved upon by increasing the mass ratio at which the orbital shrinkage decreases (i.e. where the transition occurs in Figure 10), this would lead to a decreased total number of WD+MS binaries in the final sample (Appendix B).
Compared to the E24 model, our model leads to a WD distribution with a clear deficit at high masses, in better agreement with observations. There are two contributing factors. First, our fiducial model takes into account the early termination of the core growth due to binary interaction. While this has a non-negligible effect, just accounting for binary truncation of the core’s mass growth is insufficient to match the observed WD mass distribution. More importantly, the post-MT orbital separation being a function of mass ratio not only affects the luminous star mass distribution, but also significantly reduces the fraction of systems with massive donors (with low $q_{\rm init}$) that form wide orbits, and therefore the fraction of massive WDs in the final sample.
Next, the period distribution of our model also better matches the true distribution compared to the E24 model. Since this is largely determined by the assumed orbital evolution for which we adopt more realistic boundaries of MT from an RGB vs. AGB donor compared to the E24 model (as described in Section 5.1), it suggests that the assumptions made are likely reasonable. It is particularly interesting that only a fraction of MT products from an AGB donor – those with $q_{\rm init}\gtrsim 0.4$ – end up in these relatively wide AU-scale orbits. Since the stellar radii peak during thermal pulses, we might expect interaction to occur most commonly during this phase when the envelope is most loosely bound. But if this is the relevant physical mechanism, our model requires the envelope to be efficiently ejected in only about $5-15\%$ of cases (given the uncertainties arising from the calculation of color excess; Section 3.1.1). If AGB mass transfer always resulted in a wide orbit, the model would significantly overpredict the observed number of WD+MS binaries.
Although our modeling is agnostic to whether AGB mass transfer is more akin to stable mass transfer or CEE, the mass ratio-dependent $\beta$ assumed in our model could also naturally be interpreted as a criterion for mass transfer stability: that is, orbits remain wide if the accretor-to-donor mass ratio exceeds $\sim 0.4$, and otherwise shrink significantly.
@@ 6.1 Fiducial model
By construction, the eccentricity distribution now closely matches that of the real sample. We emphasize our simulations self-consistently account for the eccentricity bias in the astrometric orbits (e.g. Wu et al., 2024), and still we require an underlying eccentricity distribution that peaks around $0.1$ to reproduce the observed eccentricities of WD+MS binaries in astrometric orbits. While some previous works have suggested that these non-zero eccentricities are a result of triple dynamics (e.g. Belloni et al., 2024), we consider it highly unlikely that all eccentric binaries – a large majority of the sample – have unseen tertiaries.

### v3

@@ 5.3 Eccentricity evolution
We alter the eccentricity evolution so that the final eccentricity, $e_{f}$, of the wide post-MT systems are related to their initial values, $e_{i}$, by $e_{f}=0.8e_{i}^{4}+0.04$. A gaussian noise with a standard deviation of 0.05 is also added, and an upper bound of $e_{f}=1$ was imposed.
This is an ad-hoc prescription which simply allows us to understand how the complicated selection function in forming the final observed sample affects its eccentricity distribution. However, simulations of CEE from Glanz & Perets (2021) found a correlation between initial and final eccentricities at the end of the dynamical phase which motivates a choice of this kind relating the two. There are also several theoretical studies of stable MT occurring in eccentric orbits, which find that orbits may not circularize as efficiently as previously assumed (Hadjidemetriou, 1969; Sepinsky et al., 2007; Hamers & Dosopoulou, 2019; Rocha et al., 2025). Implementing such models for the study of wide WD+MS binaries is a promising avenue for future work.
5.4 Models for orbital evolution under MT
We modify the conditions that lead to the formation of close and wide post-interaction binaries from those described in Section 2. We test several different models, which are described below.
Each model of the orbital evolution produces a new simulated population and a mock NSS catalog (Section 2). We process each mock catalog through the Shahaf et al. (2024) selection algorithm (Section 3) to obtain the final NCE sample of WD+MS binary candidates. Table 1 summarizes the features of each model and the resulting mock sample.
Section Description Total # $P_{\rm orb}$ [d] $_{1} [M] M$$_{\rm 2,dark} [M$]
Observed 3 Sample of Gaia WD+MS binary candidates from Shahaf et al. (2024). 3145 $640\pm 234$ $0.86\pm 0.14$ $0.61\pm 0.12$
E24 2 10% of post-AGB MT orbits remain wide at half their initial separations. All other post-MT binaries end up in 11\,d periods. 2900 $663\pm 184$ $0.80\pm 0.20$ $0.70\pm 0.21$
Stable MT 5.4.1 Use Temmink et al. (2023) quasi-adiabatic criterion for critical mass ratios. For stable MT, evolve orbits according to Soberman et al. (1997) with $\beta=1$. For systems with RGB donors, calculate WD masses so that final periods are consistent with relation from Rappaport et al. (1995). All other post-MT binaries end up in 11\,d periods. 9900 $602\pm 245$ $0.80\pm 0.20$ $0.67\pm 0.15$
$\alpha\lambda$ 5.4.2 Separations of all post-AGB MT orbits determined by the $\alpha$-formalism, with $\alpha\lambda=10^{4}$. All other post-MT binaries end up in 11\,d periods. 31000 $543\pm 252$ $0.82\pm 0.19$ $0.72\pm 0.19$
$\gamma$ 5.4.3 Separations of all post-AGB MT orbits determined by the $\gamma$-formalism, with $\gamma=2$. All other post-MT binaries end up in 11\,d periods. 17000 $554\pm 260$ $0.81\pm 0.20$ $0.74\pm 0.14$
Empirical 5.4.4 Fixed critical mass ratio of 0.38. For stable MT from AGB donors, evolve orbits according to Soberman et al. (1997) with $\beta=1$. All other post-MT binaries end up in 11\,d periods. 3100 $609\pm 234$ $0.89\pm 0.15$ $0.62\pm 0.06$
Table 1: Summary of the observed NCE sample and simulated NCE samples resulting from different models of the orbital evolution due to MT. Total numbers are rounded to 2 significant figures. The three rightmost columns list the median and standard deviation of each parameter within each sample.
5.4.1 Model for stable MT
First, we describe a model which assumes that the AU-scale WD+MS binaries are post-stable MT systems. In this case, we need to determine whether each interacting system undergoes stable or unstable MT. Temmink et al. (2023) ran a large grid of 1D binary models and determined the “quasi-adiabatic” stability boundary. The corresponding critical accretor-to-donor mass ratio above which MT remains stable, $q_{\rm qad}$, depends on the mass and radius of the donor (Figure 8 of Temmink et al. 2023).
We determine the orbital evolution due to stable MT using relations derived by Soberman et al. (1997). We assume fully non-conservative mass transfer with isotropic re-emission from the accretor (i.e. $\beta=1.0$). In this limit, the ratio of the final to initial separation, $a_{f}/a_{i}$, is given by:
$\frac{a_{f}}{a_{i}}=\left(\frac{q_{\rm f}}{q_{\rm MT}}\right)^{3}\left(\frac{1+q_{\rm f}}{1+q_{\rm MT}}\right)e^{2(1/q_{\rm f}-1/q_{\rm MT})}$ (5)
where $q_{\rm f}=M_{2,\rm init}/M_{\rm WD}$ (Soberman et al. 1997 use donor-to-accretor mass ratios, which we have inverted here). If $q_{\rm f}$ is sufficiently large compared to $q_{\rm MT}$, $a_{f}/a_{i}>1$, and the orbit widens.
Temmink et al. (2023) consider fully conservative MT. Typically, non-conservative MT is expected to lead to increased stability, so the critical mass ratios used can be regarded as upper limits. Accordingly, our modeling may be expected to under-predict the number of stable MT products. In fact, we show in Section 6.1 that the model significantly over-predicts the number of detectable systems.
Additionally, Temmink et al. (2023) do not model AGB donors with masses below $2\,M_{\odot}$, as they find that stars in this mass range do not expand significantly beyond the tip of the RGB until the start of thermal pulses, where Temmink et al. (2023) end their simulations. We find that AGB donors less massive than $2\,M_{\odot}$ make up a sizable fraction ($\sim 30\%$) of all AGB donors. By default, we assume that all of these systems undergo stable MT. For donors above $2\,M_{\odot}$, we compare the mass ratio at the onset of MT to $q_{\rm qad}$ obtained by interpolating tabulated values from Temmink et al. (2023).
It has long been appreciated that many low-mass WDs formed through stable MT from RGB donors lie on a tight relation between period and WD mass (e.g. Refsdal & Weigert, 1971; Chen et al., 2013; Istrate et al., 2014; El-Badry & Rix, 2022). Rappaport et al. (1995) derived a fitting formula for this relation which remains widely used today. Therefore, for donors that overflow their Roche lobes on the RGB, we solve for final WD masses by equating the relation from Rappaport et al. (1995) to Equation 5. While Rappaport et al. (1995) only considered donors with masses up to $2\,M_{\odot}$, this is not a problem here, as MT with more massive RGB donors become unstable in the vast majority of cases. For the few exceptions, we continue using the isolated WD IFMR from Weidemann (2000), which is our default choice unless stated otherwise.
For systems that interact with AGB donors, we also adopt the Weidemann (2000) IFMR. However, if the core growth is effectively terminated by the onset of MT, we might expect less massive WDs for a given initial stellar mass. An alternative approach would then be to interpolate over a grid of stellar radii and masses to estimate the core mass at Roche lobe overflow (analogous to how donor masses were obtained in Section 5.2). We test this in Appendix A.4 and find no significant improvement in the match between the resulting mock and observed NCE samples, so we choose not to truncate core growth in the default model.
All interacting systems which experience unstable MT (i.e. CEE) are assumed to end up in short (11\,d) periods and circular orbits. We make the same assumption for cases where MT begins before the donor is on the RGB (i.e. Hertzsprung gap/sub-giant donors). This is inconsequential to our results as such systems produce binaries containing very low-mass WDs with short orbital periods, which are unlikely to make it into the NSS catalog and make it past the AMRF cuts.
Moving forward, these assumptions will be referred to as the “stable MT model”. The key components of the model are illustrated in Figure 11.
Figure 11: Schematic summarizing the assumed orbital evolution of the zero-age binary population in the stable MT model, which leads to the formation of short-period and wide post-MT binaries. Detailed descriptions of the components can be found in Section 5.4.1.
5.4.2 An $\alpha$-formalism model
It is uncertain whether stable mass transfer or CEE is a more accurate description of the binary interaction that leads to the formation of AU-scale post-MT binaries. Therefore, we also run a model using the $\alpha$-formalism prescription (Livio & Soker, 1988; de Kool, 1990), often used to describe orbital evolution under CEE (e.g. Zorotovic et al., 2010; Davis et al., 2012; Scherbak & Fuller, 2023).
@@ 5.4.2 An $\alpha$-formalism model
Here, we test whether the $\alpha$-formalism alone can predict the properties of WD+MS binaries formed via interaction on the AGB. At its core, the formalism is a statement of energy conservation:
@@
where $E_{\rm bind}$ is the binding energy of the donor’s envelope, and the right-hand side is a fraction $\alpha$ of the orbital energy loss. $M_{i}=M_{\rm 1,MT}$ is the initial mass of the WD progenitor (i.e. the donor). In the classic formulation, $E_{\rm bind}=-(GM_{i}M_{i,\rm env})/(\lambda R_{i})$ where $M_{i,\rm env}$ and $R_{i}$ are the envelope mass and radius of the donor, and $\lambda$ is a free parameter which encodes information about the envelope structure, often set to $\sim 0.5$. Taking $R_{i}=r_{\rm L,peri}$, the Roche lobe radius at periastron, nd $M_{i,\rm env}=M_{i}-M_{\rm WD}$, we can solve for $a_{f}$ with one free parameter, namely the product $\alpha\lambda$:
$\frac{a_{f}}{a_{i}}=\frac{M_{\rm WD}}{M_{i}}\left(1+\frac{2M_{i,env}a_{i}}{M_{\star}R_{i}\alpha\lambda}\right)^{-1}$ (7)
We apply this to all systems that experience MT from AGB donors. All other systems which interact with RGB donors are assumed to end up in 11\,d circular orbits.
@@
A value of $\alpha\lambda=1$ is typical under the standard assumptions because $\alpha$ is defined to be less than or equal to 1 and considering only the gravitational binding energy, we expect $\lambda$ to be on the order unity. A higher value may be understood as the donor having a particularly loosely bound envelope, which may be the case if it is undergoing a thermal pulse on the AGB, or if additional sources of energy (e.g. recombination) aid in the unbinding (e.g. Ivanova et al., 2013, 2015; Belloni et al., 2024; Yamaguchi et al., 2024a, b). Several earlier works have found large (even negative, considering recombination) values of $\lambda$ for AGB stars (e.g. Paczyński & Ziółkowski, 1968; Han et al., 1994; Dewi & Tauris, 2000).
@@
We will refer to this model as the $\alpha\lambda$ model” hereafter. Although the $\alpha$-formalism is not expected to provide an accurate description of the mass transfer process, we include this model as it may be useful to have an effective value of $\alpha\lambda$ to predict properties of PCEBs without having to use detailed evolutionary models, which themselves struggle to predict the outcomes of CEE. The results of using this model can be found in Section 6.2 and its components are illustrated in Figure 12.
Figure 12: Components of the $\alpha\lambda$ and $\gamma$ models distinct from the stable MT model illustrated in Figure 11. Detailed descriptions are provided in Sections 5.4.2 and 5.4.3.
5.4.3 A $\gamma$-formalism model
We implement another commonly used prescription in the CEE literature, the $\gamma$-formalism. The $\gamma$-formalism was originally introduced to describe the first MT phase in the formation of double helium WD binaries (He DWDs). It is derived from angular momentum balance and may be appropriate if the common envelope is brought into co-rotation with the orbit such that there are no drag forces and thus little orbital shrinkage (Nelemans et al., 2000; Nelemans & Tout, 2005). Indeed, the orbit can widen under this formalism, and the change in separation is given by:
$\frac{a_{f}}{a_{i}}=\left(\frac{M_{i}}{M_{\rm WD}}\right)^{2}\left(\frac{M_{\rm WD}+M_{\star}}{M_{i}+M_{\star}}\right)\left(1-\gamma\frac{M_{\rm env}}{M_{i}+M_{\star}}\right)^{2}$ (8)
where the parameters are the same as in equation 7. We only apply Equation 8 for systems with AGB donors, and we adopt $\gamma=2$ as the fiducial value.
Physically, $\gamma$ may be thought of as describing the average amount of specific angular momentum carried away by mass lost from the system. Although only a narrow range of $\gamma\sim 1.5-1.7$ has been shown to be consistent with all He DWDs (Nelemans et al., 2000; Nelemans & Tout, 2005), the constraining power of this parameter has since been called into question (Webbink, 2008; Woods et al., 2011; Ivanova et al., 2013). The $\gamma$ formalism has also been interpreted to be a description of stable MT as opposed to the CEE (Ivanova et al., 2013; Scherbak & Fuller, 2023). Nevertheless, since this model remains commonly used in the literature, we apply it to the WD+MS sample ($\gamma$ model”) and present results in Section 6.3.
We also mention here the ‘grazing envelope evolution’ (GEE; Soker 2015) where steady jets from the donor launches material at the outskirts of the accretor’s envelope. This can lead to the system avoiding the CEE and forming wide, eccentric orbits (Kashi & Soker, 2018; Abu-Backer et al., 2018). While work remains to define the parameter space over which GEE is expected to occur and model the process with 3D hydrodynamical simulations, this is another possible scenario to explain the intermediate post-AGB systems and is an additional avenue to test in the future.
5.4.4 Fixed critical mass ratio
As will be described in Section 6, we find that all physically motivated models described above fail to reproduce some feature of the observed WD+MS binary sample. To better match the population demographics of the observed sample and infer the intrinsic population of post-MT binaries, we construct an “empirical model” that is tuned to produce an AMRF sample that matches the observed one.
For this model, we set a fixed value for the critical mass ratio, $q_{\rm crit}$ at the onset of MT from AGB donors. For $q>q_{\rm crit}$, we implement the Soberman et al. (1997) relation for $a_{f}/a_{i}$, as in the stable MT model (Equation 5). We ultimately choose a critical accretor-to-donor mass ratio of $q_{\rm crit}=0.38$, which leads to a final WD+MS binary sample similar in size to that observed (Section 6.4). This results in about $5\%$ of post-AGB interaction systems having orbits with final periods of $P_{\rm orb}=100-1000\,$d. Although $\sim 63\%$ of interacting systems in our model have $q>q_{\rm crit}$, the majority end up with $P_{\rm orb}>1000\,$d, and will thus only be accessible in future Gaia data releases.
In the empirical model, we assume that all binaries with RGB or Hertzsprung-gap donors donors shrink to small orbits with 11\,d periods, preventing them from entering the NSS catalog. We make this choice because in the observed sample, there is no group of systems with periods and WD masses that lie on the expected relation for formation through stable MT from RGB donors (Rappaport et al. 1995; Section 6).
It is natural to interpret the empirical model as a measurement of the average stability boundary for AGB donors. However, we emphasize that the adopted value of $q_{\rm crit}=0.38$ is purely empirical, chosen only to reproduce the observed mass and period distributions. The key input here is simply that binaries with more extreme mass ratios shrink to small orbits, whereas those with closer-to-equal mass ratios remain relatively wide. In reality, $q_{\rm crit}$ is likely to depend on the state of the donor.
@@ 6 Results of modified models
6.1 Stable MT model
We begin with results of the stable MT model, which makes use of the critical mass ratios from Temmink et al. (2023) and orbital evolution from Soberman et al. (1997) (Section 5.4.1). The resulting mock NCE sample is shown in $P_{\rm orb}-M_{\rm 2,dark}$ space in panel C of Figure 8. Distributions of several other parameters are shown in Figure 7 with an orange dashdot border.
The total number of systems in this mock sample is approximately 3 times greater than observed, suggesting that there are too many stable MT products which remain in sufficiently wide orbits for detection. This is despite the fact that the critical mass ratios from Temmink et al. (2023) assumes fully conservative MT, which is expected to represent a lower limit on the fraction of MT episodes that remain stable. While we assume fully non-conservative MT in the default model, we note that the discrepancy persists if we instead assume fully conservative MT (i.e. $\beta=0$ and $\epsilon=1$ in the notation of Soberman et al. 1997). The tension also persists even if we assume that all systems with AGB donors less massive than $2\,M_{\odot}$ experience CEE and shrink to small orbits, in which case the total number of WD+MS binaries in the final sample over-predicted by a factor of $\sim 2.5$.
The stable MT model also predicts a strip of systems hosting low-mass WDs below $\sim 0.55\,M_{\odot}$ that is not present in the observed sample. These are products of stable MT on the RGB and lie on the period-WD mass relation from Rappaport et al. (1995). The fact that these systems are not observed is another sign that the critical mass ratios from Temmink et al. (2023) may be too extreme, predicting MT to be more stable than it is in nature.
Compared to the E24 sample, there are more systems with $P_{\rm orb}\lesssim 1\,$yr, in better agreement with the observed sample. This is primarily a result of the average value of $a_{f}/a_{i}$ calculated using Equation 5 being greater than the 0.5 used in the E24 model. This means that while there are fewer systems in initially wide orbits that interact and shrink to periods short enough to be constrained in DR3, there are more systems in initially smaller orbits that remain wide enough to be detected. The adoption of more realistic boundaries for MT from RGB vs. AGB donors (Section 5.1) also affects which systems ultimately end up in detectable orbits.
The stable MT model also predicts fewer high-mass ($\gtrsim 1.0M_{\odot}$) WDs compared to the E24 model. This is because massive donors, which form massive WDs, tend to have more unequal mass ratios when MT commences, making mass transfer less likely to remain stable according to Temmink et al. (2023). This was not modeled in the E24 model, where the outcome of binary interactions is independent of the mass ratio.
Still, this model significantly overpredicts the number of WDs above $\sim 0.7M_{\odot}$ compared to the observed sample – another hint that its stability criteria may be too loose. In addition, the primary mass distribution is similar to that of the E24 model, being broader and peaked at a lower value than observed.
@@ 6.1 Stable MT model
By construction, the eccentricity distributions of our updated models all match the observed sample. We emphasize that our simulations self-consistently account for eccentricity bias in the astrometric orbits (e.g. Wu et al., 2025), and still we require an underlying eccentricity distribution that peaks around $0.1$ to reproduce the observed eccentricities of WD+MS binaries in astrometric orbits. While some previous works have suggested that these non-zero eccentricities are a result of triple dynamics (e.g. Belloni et al., 2024), we consider it unlikely that all eccentric binaries – a large majority of the sample – have unseen tertiaries.
6.2 $\alpha\lambda$ model
Next, we discuss predictions of the $\alpha\lambda$ model (Section 5.4.2), where the $\alpha$-formalism prescription is used to describe the orbital shrinkage resulting from a common envelope phase with an AGB donor.
Firstly, setting $\alpha\lambda=1$ corresponds to $f_{\rm shrink}\lesssim 0.1$. This results in very few systems entering the NSS catalog and ultimately, the final NCE sample. This issue has already been explored by several recent works (e.g. Yamaguchi et al., 2024a, b; Belloni et al., 2024) – the AU-scale WD+MS binaries discovered by Gaia cannot be explained as a result of CEE with $\alpha\lambda\sim 1$ (Zorotovic et al., 2010; Davis et al., 2012; Toonen & Nelemans, 2013; Scherbak & Fuller, 2023). Instead, they require for the donor’s envelope to be almost unbound.
In Panel D of Figure 8, we show the results of setting $\alpha\lambda=10^{4}$ on the $P_{\rm orb}-M_{\rm WD}$ space, corresponding to a range of $f_{\rm shrink}\sim 0.15-0.65$. In this case, the model predicts about 10 times more WD+MS binaries in the final sample than are observed. To remedy this tension, we can assume that only 10% of systems that interact on the AGB have $\alpha\lambda=10^{4}$, while the remainder have $\alpha\lambda\sim 1$ and form short-period PCEBs. Still, the model predicts proportionally too many systems with $P_{\rm orb}\lesssim 1\,$yr and too few with $P_{\rm orb}\gtrsim 750\,$d. This is because the intrinsic period distribution of WD+MS binaries produced by this model rises more rapidly towards shorter periods compared to the stable MT or empirical models (Figure 14). Further increasing $\alpha\lambda$ does not alleviate this problem as $a_{f}/a_{i}$ asymptotes to the ratio of the WD mass over its progenitor mass as $\alpha\lambda\rightarrow\infty$ (Equation 7). Moreover, this model does not reproduce the highly peaked WD mass distribution. This is because at sufficiently large values of $\alpha\lambda$, $a_{f}/a_{i}$ depends weakly on the initial mass ratio of the system. Given these limitations, we conclude that the $\alpha$-formalism is not well suited to describe the formation of these wide WD+MS binaries. All of the same issues exist for a more modest value of $\alpha\lambda=100$, but with an even more disproportionate abundance of systems with $P_{\rm orb}\lesssim 1\,$yr.
6.3 $\gamma$ model
We present predictions of the $\gamma$ model (Section 5.4.3), where the $\gamma$-formalism was used to predict the amount of orbital shrinkage that occurs due to MT from an AGB donor.
As shown in Panel E of Figure 8, we find that the predicted population suffers many of the same problems as the population predicted by the $\alpha\lambda$ model, though less severely. We find that $\gamma=2.0$ results in slightly more systems at the longest periods compared to the $\alpha\lambda$ model. A lower value of $\gamma=1.5$, as inferred for He DWDs, significantly overproduces the number of systems at $P_{\rm orb}\lesssim 1\,$yr. It is unsurprising that the model predicts long periods, as the $\gamma$-formalism was introduced precisely to allow for wider post-MT separations that could be not be produced by $\alpha\lambda$ model.
We see that the WD mass distribution is double peaked with a trough at $\sim 0.6\,M_{\odot}$. This is a consequence of the behavior of Equation 8, which has a non-monotonic relation of the WD mass with a local minimum around this value for $\gamma=2.0$.
Moreover, we find the primary mass distribution remains broad compared to the observed sample. This is because for values of $a_{f}/a_{i}$ such that the final separation is detectable with astrometry, there is a weak dependence on the accretor (and ultimately, the final primary) mass. Moreover, for a single value of $\gamma=2.0$, there is about five times more systems that enter the final WD+MS binary sample than observed.
6.4 Empirical model
Lastly, the distributions of stellar and orbital parameters of the mock NCE sample from our empirical model are plotted with a solid purple border on Figure 7. In Panel F of Figure 8, we also plot the systems in $P_{\rm orb}-M_{\rm WD}$ space.
Compared to the E24 and stable MT models, this model leads to a primary mass distribution that is narrower and peaked at a higher value, and a WD mass distribution with a stronger deficit above $0.75\,M_{\odot}$, in good agreement with the observed distributions. This occurs because in the empirical model, the MT stability is a step function in $q$, such that only systems with initial mass ratios above $0.38$ remain wide. This is a stricter cut than the stable MT model, where stability is maintained for a wider range of mass ratios. In this scheme, less massive donors with more massive companions are strongly favored to form wide post-MT systems that are detectable by astrometry. This also alleviates the discrepancy in the total number of systems in the final NCE sample found for the stable MT model, producing a final NCE sample with almost exactly the same number of systems as observed.
By construction, there is no strip of stable MT products from RGB donors predicted for the empirical model. This suggests that the majority of systems that interact with RGB donors undergo CEE and shrink to short orbital periods. The period distribution of the empirical model matches the observed distribution reasonably well.
Comparing the color gradient of the data points in Panels A and F of Figure 8, we note that that the correlation between the WD and MS star masses is stronger in the simulated sample than in the observed sample. A correlation between the two is expected, as systems hosting more massive MS stars must be accompanied by a more massive WD for them to make it past the AMRF cut. This leads to all systems in both the observed and simulated samples having WD-to-MS mass ratios above $\sim 0.6$. However, the spread above this lower bound is consistently narrower for the simulated sample, across all models. This is also reflected in the different slopes of the dense strip of points corresponding to the WD+MS binaries on the AMRF-$M_{1}$ space, as seen in Figure 6. This suggests that the empirical model does not capture all the physics relevant to setting the WD or primary masses, perhaps due to truncation of AGB stars’ core growth after MT begins.